### v3

@@ 4.4 Results of RQ3: Supported Architecting Activities
Datasets and sources in AS - Text-based dataset (e.g., [S15][S79]) is the frequently used dataset in AS, whereas, Q&A sites, Wiki, issue tracking systems, and technical blogs and tutorials are the main sources in this activity.
@@
Architecture Conformance Checking (ACC) is conducted to check and evaluate if the implemented architecture conforms to the intended architecture, and this activity can help architects identify and correct architectural violations and further avoid constant architecture erosion during the software development life cycle (Pruijt et al., 2013). 9.6% (10 out of 104) of the studies investigated ACC activity.
@@
Datasets and sources in ACC - Code-based (e.g., [S85][S81]) and text-based (e.g., [S51]) datasets are the main datasets used in ACC. Version control systems, issue tracking systems, search engines, and Wiki are the common sources for collecting data in ACC (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)).
Finding 5: Eleven architecting activities can be supported by the mined architectural information, in which architecture understanding (44.2%, 46 out of 104)), architecture maintenance and evolution (40.4%, 42 out of 104)), and architecture recovery (32.7%, 34 out of 104 studies) are the top three most supported architecting activities. Moreover, architectural description is the most common type of mined architectural information that supports all the architecting activities. Finding 6: Various types of problems (e.g., recovering architectural description via mining architectural views and models) have been addressed in different architecting activities (e.g., ARec). Finding 7: Classification and clustering-based approaches are the most frequently employed approaches for mining architectural information to address various problems in different architecting activities.
@@ 4.5.1 Approaches
The selected studies have proposed and used different approaches to mine architectural information from different sources. We analyzed and synthesized the approaches used for mining architectural information by following the coding steps for data synthesis as described in Section 3.2.5. We observed that some selected studies did name their approaches used to mine architectural information and others did not. If a study does not provide a name for the proposed approach, we labeled it as “Unnamed” (see the Automatic Approaches sheet in the dataset (de Dieu et al., 2023)). Moreover, we found that the proposed approaches have different levels of automation (e.g., automatic, semi-automatic). In addition, these approaches have been applied in different types of tasks (e.g., classification, clustering) for mining architectural information from various sources. Thus, in this SMS, we utilized a two-level categorization to categorize the approaches for mining architectural information. At the first level, we divided the approaches in three high-level categories based on the level of automation: (1) automatic, (2) semi-automatic, and (3) manual. At the second level, we further categorized the approaches in three categories based on different types of tasks in mining architectural information: (1) architectural information classification, (2) architectural information clustering, and (3) architectural information retrieval. In total, we gathered 95 approaches from the selected studies of which 60% (i.e., 57 out of 95) are automatic approaches (see the Automatic Approaches sheet in the dataset (de Dieu et al., 2023)), 34.7% (i.e., 33 out of 95) are semi-automatic approaches (see the Semi-automatic Approaches sheet in the dataset (de Dieu et al., 2023)), and 5.3% (5 out of 95) are manual approaches (see the Manual Approaches sheet in the dataset (de Dieu et al., 2023)). Architectural information classification-based approaches count 52.6% (50 out of 95), followed by clustering-based approaches (26.3%, 25 out of 95), and information retrieval-based approaches (15.8%, 15 out of 95). Furthermore, to help and ease the applicability of these approaches in practice, we also present the sources that the proposed approaches can be applied on.
@@
Architectural information classification refers to the approaches that mine architectural information based on classification techniques. Architectural information classification-based approaches are the most frequently 56.2% (50 out of 95) used approaches in mining architectural information. Classification techniques learn from the data input provided to them and then use this learning to classify new observations (Kotsiantis et al., 2006). There are two types of architectural information classification-based approaches that were employed in the selected studies: binary classification-based approaches (e.g., [S6][S48][S59]) and multi-class classification-based approaches (e.g., [S40][S76][S79]). Architectural information classification-based approaches are applied either automatically, semi-automatically, or manually (see the Automatic Approaches, Semi-automatic Approaches, and Manual Approaches sheets in the dataset (de Dieu et al., 2023)). For example, in [S40], Mondal et al. proposed an approach named Archinet for mining and classifying architectural changes from code and commit messages in GitHub. Specifically, they formulated this problem as a multi-class classification task that automatically classifies architectural changes into four categories. In [S6], Bi et al. developed a semi-automatic dictionary-based mining approach to extract QAs and Architectural Tactic (AT) related discussions in Stack Overflow (SO) posts. They formulated this problem as a binary classification task, which automatically identifies QA-AT related discussions from SO posts. Then the authors went on to manually structure the design relationships between QAs and ATs used in practice, and build a knowledge base of how developers use ATs with respect to QA concerns. The proposed approach achieved good performance and F-measure. In [S6], Malavolta et al. employed a manual approach (i.e., qualitative analysis) to mine and elicit evidence-based architectural guidelines for open-source ROS-based software from GitHub, BitBucket, and GitLab. Their qualitative data analysis yielded 39 guidelines for architecting robotics software, which can be used by roboticists to architect their systems to achieve particular quality requirements.
@@
Architectural information clustering denotes the approaches that employ clustering-based techniques to mine architectural information from various sources. Architectural information clustering-based approaches count 26.3% (25 out of 95 approaches). The clustering techniques aim at reducing the amount of data by categorizing or grouping similar data items together into subsets or clusters (Saxena et al., 2017). The goal is to create clusters with internal coherence, placing similar objects in the same group, and assigning dissimilar objects to different groups. In this sense, documents that belong to a certain cluster should be as similar as possible and dissimilar from documents in other clusters. We identified two types of architectural information clustering-based approaches employed in the selected studies: partitional clustering (e.g., [S43]) and hierarchical clustering (e.g., [S68]). In addition, these approaches operate in two ways: automatic and semi-automatic (see the Automatic Approaches and Semi-automatic Approaches sheets in the dataset (de Dieu et al., 2023)).
@@ 4.5.2 Tools
General tools are provided to analyze, process, and mine various information, not limited to architectural information. For example, Stanford Parser is a general NLP tool frequently used in textual information prepossessing tasks, and Gokyer et al. [S66] used Stanford Parser for analyzing the grammar of textual architecture information in Wiki. We provide the collected general tools, their descriptions, URL links, and the relevant studies in which these general tools were used in the General Tools sheet in the dataset (de Dieu et al., 2023).
Dedicated tools are developed for specifically mining architectural information. We found that dedicated tools have two levels of automation (i.e., automatically and semi-automatically) and have been applied in different types of tasks (e.g., clustering) for mining architectural information from various sources. Thus, similar to what we did when analyzing architectural information mining approaches, we utilized a two-level categorization to categorize the dedicated tools for mining architectural information. At the first level, we divided these tools in three high-level categories based on the level of automation: (1) automatic (82%, 32 out of 39 tools) and (2) semi-automatic (17.9%, 7 out of 39 tools) (see the Automatic Tools and Semi-automatic Tools sheets in the dataset (de Dieu et al., 2023)). At the second level, we further categorized the dedicated tools in three categories based on different types of tasks in mining architectural information: (1) architectural information classification (35.9%, 14 out of 39 tools), (2) architectural information clustering (38.5%, 15 out of 39 tools), and (3) architectural information retrieval (25.6%, 10 out of 39 tools). Moreover, to help and ease the applicability of these approaches in practice, we also present the sources that the proposed approaches can be applied on. The detail information about the dedicated tools is shown in the Automatic Tools and Semi-automatic Tools sheets in the dataset (de Dieu et al., 2023). Specifically, these sheets list the categories of different tasks in mining architectural information, tool names, sources used, mined architectural information, URL links, and the relevant studies in which these dedicated tools were proposed. Note that, the results in the “URL link” column were collected from the selected studies. Moreover, it is worth noting that some URL links of the tools were not accessible when we conducted this SMS. Therefore we provide the links of these tools with a note “Not accessible”. On the other hand “Not provided” means that the selected study does not provide a URL link to the proposed tool.
@@
Finding 8: We gathered 95 approaches for mining architectural information, in which 60% (i.e., 57 out of 95) are automatic approaches, 34.7% (i.e., 33 out of 95) are semi-automatic approaches, and 5.3% (5 out of 95) are manual approaches. Architectural information classification is the most common task supported by the approaches 52.6% (50 out of 95) in mining architectural information. Finding 9: We collected 56 tools for mining architectural information, in which 30.4% (17 out of 56) are general tools and 69.6% (39 out of 56) are dedicated tools. Weka and Stanford POS Tagger are the most common used general tools in mining architectural information.
@@ 4.6 Results of RQ5: Challenges
The challenges refer to the obstacles or issues encountered in mining architectural information. As mentioned in Section 3.2.5, we used open coding & constant comparison and descriptive statistics to analyze the extracted data item (i.e., D10, see Table 4) and identified the challenges faced in mining architectural information. Our data synthesis yielded four categories, of which one was categorized as “Others” (referring to codes that do not fit into the already generated subcategories), and seven subcategories. Table 6 shows these categories and subcategories, their descriptions along with relevant studies.
@@
Architectural element description (26)
Vagueness or ambiguity in architectural element description
[S1] [S15] [S11] [S18] [S20] [S40] [S42] [S45] [S63] [S72] [S77]
11
Incompleteness in architectural element description
[S1] [S6] [S10] [S15] [S40] [S48] [S58] [S63] [S72]
9
Redundancy in architectural element description
[S40] [S45] [S58] [S63] [S72] [S83]
6
Approach and tool (23)
Limitation of approaches and tools
[S27] [S30] [S59] [S64] [S88] [S89] [S90] [S91] [S94] [S97] [S100] [S101] [S102] [S104]
14
Applicability of approaches to heterogeneous architectural data
[S17] [S24] [S25] [S27] [S32] [S48] [S64]
7
Feature selection
[S5][S20]
2
Architectural dataset (17)
Size of architectural dataset for ML model training
[S4] [S13] [S15] [S25] [S40] [S45] [S50] [S55] [S64] [S72]
10
Quality of architectural dataset for ML model training
[S1] [S6] [S13] [S14] [S15] [S24] [S40]
7
Others (10)
[S10] [S13] [S14] [S28] [S34] [S43] [S46] [S48] [S55] [S59]
10
Architectural element description: As shown in Figure 7 (the results of RQ2 in Section 4.3), the selected studies used diverse sources for mining architectural information. However, the description of architectural elements in these sources is critical for architectural information mining approaches and tools. 25% (26 out of 104) of the studies reported the challenges related to architectural element descriptions encountered when mining architectural information. We further categorized these challenges into three subcategories (see Table 6).
@@
Architectural dataset is a critical asset for several architectural information mining approaches and tools to extract real and significant architectural information from the datasets. However, there are no public and suitable datasets for certain architecture design tasks. Architectural dataset is reported as a challenge in 16.3% (17 out of 104) of the studies. We further categorized the challenges in the architectural dataset category into two subcategories (see Table 6).
@@
Quality of the architectural dataset for ML model training: The architectural datasets vary much in quality, including bias, noise, imbalance, and mislabelling. There are many factors that may impact the quality of architectural datasets, such as the source of the dataset, whether the dataset has been preprocessed, and the ground truth in the dataset (Yang et al., 2020), and any of these factors will influence the performance of the approaches or tools used. Quality of the architectural dataset is the challenge of 6.7% (7 out of 104) of the selected studies. In [S24], López et al. proposed an approach named Toeska Rationale Extraction (TREx) for extracting, recovering, and exploring architectural rationale information from various sources, such as Wiki, meeting notes, emails. However, the authors stated that the effectiveness (e.g., accuracy) of the approach on extracting data (e.g., architectural rationale) is heavily dependent on the quality of the available textual architectural artifacts.
@@
Finding 10: 4 types and 8 subtypes of challenges were identified in mining architectural information. The challenges in mining architectural information are mainly derived from the descriptions of architectural elements, wherein vagueness or ambiguity in architectural element description is a prevalent challenge. Finding 11: The studies report the challenges related to the limitation of architectural information mining approaches and tools. For example, There are still lacks of effective architectural information mining approaches and tools that are applicable to various architectural data from diverse sources in mining architectural information. Finding 12: The number of quality datasets with an appropriate size available for architectural information mining approaches is also regarded a serious challenge.
@@ 5 Discussion
In this section, we revisit the findings of this SMS by interpreting the results in Section 5.1 and discussing their implications for researchers and practitioners in Section 5.2
@@ 5.1.1 Mined Architectural Information Could Support Architecting Activities
The results of RQ1 (see Section 4.2) reveal that various categories and subcategories of architectural information have been mined in software development, such as architectural description, architectural solution, and system requirement (see Table 5). Moreover, the results show that architectural description is the most mined category of architectural information. The potential reason could be that architectural description plays as a blueprint for system development, specifically, a blueprint for effectively and successfully developing large and complex systems (Clements et al., 2003). In addition, the information in architectural description, such as architectural views, which describe systems in multiple views (e.g., process view and development view), enables the architecture to be communicated to, and understood by, different stakeholders. Also, it allows stakeholders to verify whether the architecture will address their concerns (ISO/IEC/IEEE, 2011). The core benefits for mining architectural description are lowering architectural knowledge vaporization and speeding up maintenance of software systems (Borrego et al., 2019). These benefits can be achieved by capturing, sharing, and reusing architectural information (such as architectural views, architectural models, and architectural rationale) in the architectural description, thus reducing the time that new developers need to familiarize with an existing system and potentially speeding up the development.
Architectural solution is the second most mined category of architectural information. The reason could be that researchers and practitioners consider architectural solutions, such as architectural patterns (e.g., service-oriented architecture) and tactics (e.g., resource pooling) as the fundamental building blocks in the architecture design of a system, and they are used to address ASRs (Bass et al., 2012). Contrarily to changing implementation (e.g., low-level source code), once an architecture solution (e.g., an architecture pattern) is adopted and implemented, it is quite difficult and costly to change it (Bass et al., 2012). Mining architectural solutions promises significant practical benefits including increased reuse of architectures, lower costs of architectural change, and less erosion of architectural design. These benefits are obtained by offering developers access to the architectural solutions that lead to the architecture, rather than only the resulting outcomes and artifacts from the architectural solutions (Ramachandran, 2002).
The mined architectural information could support a variety of architecting activities (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)). Each architecting activity may require specific architectural information for it to be carried out effectively. For example, during architecture evaluation, architectural information, such as the benefits and drawbacks of certain architectural solutions (e.g., architectural patterns and tactics) can help architects evaluate the architectural solutions (e.g., choosing suitable architectural patterns according to the positively and negatively affected quality attributes) in specific application domains. The results of RQ3 show that eleven architecting activities can be supported by the mined architectural information, where architecture understanding is the most supported activity (i.e., 44.2% (46 out of 104)). Some mined architectural information to support AU include architectural description, architectural decision, system requirement, and architectural solution (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)). An architecture of a system should be well understood before any corrections or changes can be applied to it. AU is a prerequisite for and supportive to other architecting activities (e.g., AA, AI, AME). Architectural elements, such as architectural decisions, are vital in achieving the desired software quality, and a good understanding of these architectural elements and their relationships is important in supporting AU (Stevanetic and Zdun, 2014). Understanding architectural decisions that make an architecture is also a key factor for maintaining the software system (Shahin et al., 2014). For example, AU enables maintainers to gain a comprehensive overview of the complex dependency relationships between architectural components and connectors.
@@
Architecture maintenance and evolution is the second most supported activity. It is an established fact that good software architecture and design leads to better evolvability, maintainability, availability, software cost reduction, and so on (Bass et al., 2003). However, during development, developers tend to willingly or unwillingly deviate from the architectural patterns through compromising architecture and design process by poor or hasted architectural design decisions, which lead the architecture to different architectural problems or anomalies (e.g., architectural smells). As the system evolves, the accumulation of such problems can cause the implemented architecture to deviate away from the intended architecture, i.e., a typical case of architecture erosion (Li et al., 2022). An architecture with such problems can aggravate the brittleness of the system (Perry and Wolf, 1992), decrease architecture sustainability (Koziolek et al., 2013), and make it difficult for developers to understand the internal structure of the system (Perry and Wolf, 1992). Therefore, these facts motivate researchers and practitioners to turn their attention to developing architectural information mining approaches and tools to assist AME activity.
@@
Furthermore, sometimes architectural information is tacit and not documented at all (Ding et al., 2014). With the increase in the size and complexity of the systems, architectural information management becomes more challenging, and stakeholders need to find the right information and recover it efficiently. Searching and recovering relevant architectural information from various artifacts is a challenging task. With a large volume of artifacts, applying mining approaches can greatly facilitate the task of recovering desired architectural information. Thus, it is not surprising that architecture recovery is the third (i.e., 32.7% (34 out of 104)) most supported architecting activity (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)). Architecture can be discovered from various artifacts, such as source code and software documentation (Ducasse and Pollet, 2009). During our data synthesis, we found that most of the studies supporting AR recovered architectural information, such as architectural description (e.g., architectural views and models) from source code (e.g., [S39][S47]).
@@
The relationships between the mined architectural information (results of RQ1), sources (RQ2), and supported architecting activities (results of RQ3): We show the relationships between mined architectural information, sources, and supported architecting activities through the results of RQ1 (see Section 4.2), RQ2 (see Section 4.3), and RQ3 (see Section 4.4) (see the Architecting Activities sheet in the dataset (de Dieu et al., 2023)). This information can, for example, help researchers and practitioners know which architecting activity (e.g., architecture maintenance and evolution) can be supported by what types of mined architectural information (e.g., architectural technical debt) during the development. In addition, such information can also help practitioners know which sources store certain architectural information or elements (e.g., reusable architectural elements) for certain architecting activities (e.g., architecture reuse).

### v4

@@ 4.4 Results of RQ3: Supported Architecting Activities
Datasets and sources in AS - Text-based dataset (e.g., [S15][S79]) is the frequently used dataset in AS, whereas, Q&A sites, Wiki, issue tracking systems, and technical blogs & tutorials are the main sources in this activity.
@@
Architecture Conformance Checking (ACC) is conducted to check and evaluate if the implemented architecture conforms to the intended architecture, and this activity can help architects identify and correct architectural violations and further avoid constant architecture erosion during the software development life cycle (Rocha et al., 2017). 9.6% (10 out of 104) of the studies investigated ACC activity.
@@
Datasets and sources in ACC - Code-based (e.g., [S85][S81]) and text-based (e.g., [S51]) datasets are the main datasets used in ACC. Version control systems, issue tracking systems, search engines, and Wiki are the common sources for collecting data in ACC (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)).
Key Findings of RQ3 Finding 3: Considerable effort has been devoted to mining architectural information from software repositories to support various architecting activities (i.e., 11 architecting activities). As a result, the current research on mining architectural information can help practitioners be aware of which architecting activities (such as AU and AME) can be supported by the mined architectural information. Finding 4: Although various types of problems (e.g., recovering architectural description via mining architectural views and models) have been addressed in different architecting activities (e.g., ARec), there are still unresolved problems in certain architecting activities, including mining and recommending the implementation of certain architectural elements, such as architectural pattern implementations to assist AIm activity.
@@ 4.5.1 Approaches
The selected studies have proposed and used different approaches to mine architectural information from different sources. We analyzed and synthesized the approaches used for mining architectural information by following the coding steps for data synthesis as described in Section 3.2.5. We observed that some selected studies did name their approaches used to mine architectural information and others did not. If a study does not provide a name for the proposed approach, we labeled it as “Unnamed” (see the Automatic Approaches sheet in the Supplementary Material (de Dieu et al., 2024)). Moreover, we found that the proposed approaches have different levels of automation (e.g., automatic, semi-automatic). In addition, these approaches have been applied in different types of tasks (e.g., classification, clustering) for mining architectural information from various sources. Thus, in this SMS, we utilized a two-level categorization to categorize the approaches for mining architectural information. At the first level, we divided the approaches in three high-level categories based on the level of automation: (1) automatic, (2) semi-automatic, and (3) manual. At the second level, we further categorized the approaches in three categories based on different types of tasks in mining architectural information: (1) architectural information classification, (2) architectural information clustering, and (3) architectural information retrieval. In total, we gathered 95 approaches from the selected studies of which 60% (i.e., 57 out of 95) are automatic approaches (see the Automatic Approaches sheet in the Supplementary Material (de Dieu et al., 2024)), 34.7% (i.e., 33 out of 95) are semi-automatic approaches (see the Semi-automatic Approaches sheet in the Supplementary Material (de Dieu et al., 2024)), and 5.3% (5 out of 95) are manual approaches (see the Manual Approaches sheet in the Supplementary Material (de Dieu et al., 2024)). Architectural information classification-based approaches count 52.6% (50 out of 95), followed by clustering-based approaches (26.3%, 25 out of 95), and information retrieval-based approaches (15.8%, 15 out of 95). Furthermore, to help and ease the applicability of these approaches in practice, we also present the sources that the proposed approaches can be applied on.
@@
Architectural information classification refers to the approaches that mine architectural information based on classification techniques. Architectural information classification-based approaches are the most frequently 56.2% (50 out of 95) used approaches in mining architectural information. Classification techniques learn from the data input provided to them and then use this learning to classify new observations (Kotsiantis et al., 2006). There are two types of architectural information classification-based approaches that were employed in the selected studies: binary classification-based approaches (e.g., [S6][S48][S59]) and multi-class classification-based approaches (e.g., [S40][S76][S79]). Architectural information classification-based approaches are applied either automatically, semi-automatically, or manually (see the Automatic Approaches, Semi-automatic Approaches, and Manual Approaches sheets in the Supplementary Material (de Dieu et al., 2024)). For example, in [S40], Mondal et al. proposed an approach named Archinet for mining and classifying architectural changes from code and commit messages in GitHub. Specifically, they formulated this problem as a multi-class classification task that automatically classifies architectural changes into four categories. In [S6], Bi et al. developed a semi-automatic dictionary-based mining approach to extract QAs and Architectural Tactic (AT) related discussions in Stack Overflow (SO) posts. They formulated this problem as a binary classification task, which automatically identifies QA-AT related discussions from SO posts. Then the authors went on to manually structure the design relationships between QAs and ATs used in practice, and build a knowledge base of how developers use ATs with respect to QA concerns. The proposed approach achieved good performance and F-measure. In [S6], Malavolta et al. employed a manual approach (i.e., qualitative analysis) to mine and elicit evidence-based architectural guidelines for open-source ROS-based software from GitHub, BitBucket, and GitLab. Their qualitative data analysis yielded 39 guidelines for architecting robotics software, which can be used by roboticists to architect their systems to achieve particular quality requirements.
@@
Architectural information clustering denotes the approaches that employ clustering-based techniques to mine architectural information from various sources. Architectural information clustering-based approaches count 26.3% (25 out of 95 approaches). The clustering techniques aim at reducing the amount of data by categorizing or grouping similar data items together into subsets or clusters (Saxena et al., 2017). The goal is to create clusters with internal coherence, placing similar objects in the same group, and assigning dissimilar objects to different groups. In this sense, documents that belong to a certain cluster should be as similar as possible and dissimilar from documents in other clusters. We identified two types of architectural information clustering-based approaches employed in the selected studies: partitional clustering (e.g., [S43]) and hierarchical clustering (e.g., [S68]). In addition, these approaches operate in two ways: automatic and semi-automatic (see the Automatic Approaches and Semi-automatic Approaches sheets in the Supplementary Material (de Dieu et al., 2024)).
@@ 4.5.2 Tools
General tools are provided to analyze, process, and mine various information, not limited to architectural information. For example, Stanford Parser is a general NLP tool frequently used in textual information prepossessing tasks, and Gokyer et al. [S66] used Stanford Parser for analyzing the grammar of textual architecture information in Wiki. We provide the collected general tools, their descriptions, URL links, and the relevant studies in which these general tools were used in the General Tools sheet in the Supplementary Material (de Dieu et al., 2024).
Dedicated tools are developed for specifically mining architectural information. We found that dedicated tools have two levels of automation (i.e., automatically and semi-automatically) and have been applied in different types of tasks (e.g., clustering) for mining architectural information from various sources. Thus, similar to what we did when analyzing architectural information mining approaches, we utilized a two-level categorization to categorize the dedicated tools for mining architectural information. At the first level, we divided these tools in three high-level categories based on the level of automation: (1) automatic (82%, 32 out of 39 tools) and (2) semi-automatic (17.9%, 7 out of 39 tools) (see the Automatic Tools and Semi-automatic Tools sheets in the Supplementary Material (de Dieu et al., 2024)). At the second level, we further categorized the dedicated tools in three categories based on different types of tasks in mining architectural information: (1) architectural information classification (35.9%, 14 out of 39 tools), (2) architectural information clustering (38.5%, 15 out of 39 tools), and (3) architectural information retrieval (25.6%, 10 out of 39 tools). Moreover, to help and ease the applicability of these approaches in practice, we also present the sources that the proposed approaches can be applied on. The detail information about the dedicated tools is shown in the Automatic Tools and Semi-automatic Tools sheets in the Supplementary Material (de Dieu et al., 2024). Specifically, these sheets list the categories of different tasks in mining architectural information, tool names, sources used, mined architectural information, URL links, and the relevant studies in which these dedicated tools were proposed. Note that, the results in the “URL link” column were collected from the selected studies. Moreover, it is worth noting that some URL links of the tools were not accessible when we conducted this SMS. Therefore we provide the links of these tools with a note “Not accessible”. On the other hand “Not provided” means that the selected study does not provide a URL link to the proposed tool.
@@
Key Findings of RQ4 Finding 5: The primary studies employed three types of approaches (automatic, semi-automatic, and manual approaches) to mine architectural information from software repositories. Moreover, since manual approaches are time-consuming and require substantial human efforts, most primary studies motivated their mining approaches to reduce the manual effort in mining architectural information, where 60.0% (i.e., 57 out of 95) are automatic approaches and 34.7% (i.e., 33 out of 95) are semi-automatic approaches, and these approaches were used to solve various tasks, such as architectural information classification and clustering. Finding 6: Although various architectural information approaches (i.e., 95 approaches) have been proposed for mining architectural information from software repositories, we found that still more than half of the approaches (i.e., 56 approaches) are not supported by tools.
@@ 4.6 Results of RQ5: Challenges
The challenges refer to the obstacles or issues encountered in mining architectural information. Our data synthesis yielded four categories, of which one was categorized as “Others” (referring to codes that do not fit into the already generated subcategories), and seven subcategories. Table LABEL:Challenges shows these categories and subcategories, their descriptions along with relevant studies.
@@
Architectural element description (26) Vagueness or ambiguity in architectural element description [S1] [S15] [S11] [S18] [S20] [S40] [S42] [S45] [S63] [S72] [S77] 11
Incompleteness in architectural element description [S1] [S6] [S10] [S15] [S40] [S48] [S58] [S63] [S72] 9
Redundancy in architectural element description [S40] [S45] [S58] [S63] [S72] [S83] 6
Approach and tool (23) Limitation of approaches and tools [S27] [S30] [S59] [S64] [S88] [S89] [S90] [S91] [S94] [S97] [S100] [S101] [S102] [S104] 14
Applicability of approaches to heterogeneous architectural data [S17] [S24] [S25] [S27] [S32] [S48] [S64] 7
Feature selection [S5][S20] 2
Architectural dataset (17) Size of architectural dataset for ML model training [S4] [S13] [S15] [S25] [S40] [S45] [S50] [S55] [S64] [S72] 10
Quality of architectural dataset for ML model training [S1] [S6] [S13] [S14] [S15] [S24] [S40] 7
Others (10) [S10] [S13] [S14] [S28] [S34] [S43] [S46] [S48] [S55] [S59] 10
Architectural element description: As shown in Figure 7 (the results of RQ2 in Section 4.3), the selected studies used diverse sources for mining architectural information. However, the description of architectural elements in these sources is critical for architectural information mining approaches and tools. 25% (26 out of 104) of the studies reported the challenges related to architectural element descriptions encountered when mining architectural information. We further categorized these challenges into three subcategories (see Table LABEL:Challenges).
@@
Architectural dataset is a critical asset for several architectural information mining approaches and tools to extract real and significant architectural information from the datasets. However, there are no public and suitable datasets for certain architecture design tasks. Architectural dataset is reported as a challenge in 16.3% (17 out of 104) of the studies. We further categorized the challenges in the architectural dataset category into two subcategories (see Table LABEL:Challenges).
@@
Quality of the architectural dataset for ML model training: The architectural datasets vary much in quality, including bias, noise, imbalance, and mislabelling. There are many factors that may impact the quality of architectural datasets, such as the source of the dataset, whether the dataset has been preprocessed, and the ground truth in the dataset (Yang et al., 2023), and any of these factors will influence the performance of the approaches or tools used. Quality of the architectural dataset is the challenge of 6.7% (7 out of 104) of the selected studies. In [S24], López et al. proposed an approach named Toeska Rationale Extraction (TREx) for extracting, recovering, and exploring architectural rationale information from various sources, such as Wiki, meeting notes, emails. However, the authors stated that the effectiveness (e.g., accuracy) of the approach on extracting data (e.g., architectural rationale) is heavily dependent on the quality of the available textual architectural artifacts.
@@
Key Findings of RQ5 Finding 7: The challenges in mining architectural information are mainly derived from the descriptions of architectural elements, wherein vagueness or ambiguity in architectural element description is a prevalent challenge. Finding 8: The primary studies also report the challenges related to the limitation of architectural information mining approaches and tools. For example, there is still a lack of effective architectural information mining approaches and tools that are applicable to various architectural datasets from diverse sources in mining architectural information. Moreover, the number of quality datasets with an appropriate size available for architectural information mining approaches is also a serious challenge.
@@ 5.1.1 Mined Architectural Information Could Support Architecting Activities
The results of RQ1 (see Section 4.2) reveal that various categories and subcategories of architectural information have been mined in software development, such as architectural description, architectural solution, and system requirement (see Table LABEL:minedArchitecturalInfo). Moreover, the results show that architectural description is the most mined category of architectural information. The potential reason could be that architectural description plays as a blueprint for system development, specifically a blueprint for effectively and successfully developing large and complex systems (Clements et al., 2003). In addition, the information in architectural description, such as architectural views, which describe systems in multiple views (e.g., process view and development view), enables the architecture to be communicated to, and understood by, different stakeholders. Also, it allows stakeholders to verify whether the architecture will address their concerns (ISO/IEC/IEEE, 2011). The core benefits of mining architectural description are lowering architectural knowledge vaporization and speeding up maintenance of software systems (Borrego et al., 2019). These benefits can be achieved by capturing, sharing, and reusing architectural information (such as architectural views, architectural models, and architectural rationale) in the architectural description, thus reducing the time that new developers need to familiarize themselves with an existing system and potentially speeding up the development. Architectural solution is the second most mined category of architectural information. The reason could be that researchers and practitioners consider architectural solutions, such as architectural patterns (e.g., service-oriented architecture) and tactics (e.g., resource pooling) as the fundamental building blocks in the architecture design of a system, and they are used to address ASRs (Bass et al., 2012). Contrarily to changing implementation (e.g., low-level source code), once an architecture solution (e.g., an architecture pattern) is adopted and implemented, it is quite difficult and costly to change it (Bass et al., 2012). Mining architectural solutions promises significant practical benefits including increased reuse of architectures, lower costs of architectural change, and less erosion of architectural design. These benefits are obtained by offering developers access to the architectural solutions that lead to the architecture, rather than only the resulting outcomes and artifacts from the architectural solutions (Kruchten, 2004).
@@
The results of RQ3 show that eleven architecting activities can be supported by the mined architectural information, where architecture understanding is the most supported activity (i.e., 44.2% (46 out of 104)). Some mined architectural information to support AU include architectural description, architectural decision, system requirement, and architectural solution (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)). An architecture of a system should be well understood before any corrections or changes can be applied to it. AU is a prerequisite for and supportive to other architecting activities (e.g., AA, AI, AME). Architectural elements, such as architectural decisions, are vital in achieving the desired software quality, and a good understanding of these architectural elements and their relationships is important in supporting AU (Stevanetic and Zdun, 2014). Understanding architectural decisions that make an architecture is also a key factor for maintaining the software system (Shahin et al., 2014). For example, AU enables maintainers to gain a comprehensive overview of the complex dependency relationships between architectural components and connectors. Thus, these facts tend to motivate researchers and mine several types of architectural information from software repositories in order to facilitate AU during the architecting process. Architecture maintenance and evolution is the second most supported activity. It is an established fact that good software architecture and design leads to better evolvability, maintainability, availability, software cost reduction, and so on (Bass et al., 2012). However, during development, developers tend to willingly or unwillingly deviate from the architectural patterns through compromising architecture and design process by poor or hasted architectural design decisions, which lead the architecture to different architectural problems or anomalies (e.g., architectural smells). As the system evolves, the accumulation of such problems can cause the implemented architecture to deviate away from the intended architecture, i.e., a typical case of architecture erosion (Li et al., 2022). An architecture with such problems can aggravate the brittleness of the system (Perry and Wolf, 1992), decrease architecture sustainability (Koziolek et al., 2013), and make it difficult for developers to understand the internal structure of the system (Perry and Wolf, 1992). Therefore, these facts motivate researchers to turn their attention to developing architectural information mining approaches and tools to assist AME activity.
@@
Furthermore, sometimes architectural information is tacit and not documented at all (Ding et al., 2014). With the increase in the size and complexity of the systems, architectural information management becomes more challenging, and stakeholders need to find the right information and recover it efficiently. Searching and recovering relevant architectural information from various artifacts is a challenging task. With a large volume of artifacts, applying mining approaches can greatly facilitate the task of recovering desired architectural information. Thus, it is not surprising that architecture recovery is the third (i.e., 32.7% (34 out of 104)) most supported architecting activity (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)). Architecture can be discovered from various artifacts, such as source code and software documentation (Ducasse and Pollet, 2009). During our data synthesis, we found that most of the studies supporting AR recovered architectural information, such as architectural description (e.g., architectural views and models) from source code (e.g., [S39][S47]).
@@
The relationships between the mined architectural information (results of RQ1), sources (RQ2), and supported architecting activities (results of RQ3): We show the relationships between mined architectural information, sources, and supported architecting activities through the results of RQ1 (see Section 4.2), RQ2 (see Section 4.3), and RQ3 (see Section 4.4) (see the Architecting Activities sheet in the Supplementary Material (de Dieu et al., 2024)). This information can, for example, help researchers and practitioners know which architecting activity (e.g., architecture maintenance and evolution) can be supported by what types of mined architectural information (e.g., architectural technical debt) during the development. In addition, such information can also help practitioners know which sources store certain architectural information or elements (e.g., reusable architectural elements) for certain architecting activities (e.g., architecture reuse).